# Improving Probabilistic Diffusion Models With Optimal Covariance Matching

### v2

@@ Abstract
The probabilistic diffusion model has become highly effective across various domains. Typically, sampling from a diffusion model involves using a denoising distribution characterized by a Gaussian with a learned mean and either fixed or learned covariances. In this paper, we leverage the recently proposed covariance moment matching technique and introduce a novel method for learning the diagonal covariances. Unlike traditional data-driven covariance approximation approaches, our method involves directly regressing the optimal analytic covariance using a new, unbiased objective named Optimal Covariance Matching (OCM). This approach can significantly reduce the approximation error in covariance prediction. We demonstrate how our method can substantially enhance the sampling efficiency, recall rate and likelihood of both diffusion models and latent diffusion models.
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019) have achieved remarkable success in modeling complex real-world data across various domains (Rombach et al., 2022; Li et al., 2022; Poole et al., 2022; Ho et al., 2022; Hoogeboom et al., 2022; Liu et al., 2023). A conventional diffusion model operates in two stages: a forward noising process, indexed by $t\in[1,T]$, which progressively corrupts the data distribution into a Gaussian distribution via Gaussian convolution, and a reverse denoising process, which generates images by gradually transforming Gaussian noise back into coherent data samples. In traditional diffusion models, the generation process typically predicts only the mean of the denoising distribution while using a fixed, pre-defined variance (Ho et al., 2020). This approach often requires a very large number of steps, $T$, to produce high-quality, diverse samples or to achieve reasonable model likelihoods, leading to inefficiencies during inference.
To address this inefficiency, several works have proposed estimating the posterior covariance instead of relying on pre-defined variance values. For instance, Bao et al. (2022b) introduces an analytical form of isotropic, state-independent covariance that can be estimated from the data. This analytical solution achieves an optimal form under the constraints of isotropy and state independence. A more flexible approach involves learning a state-dependent diagonal covariance. Nichol & Dhariwal (2021) propose learning this form by optimizing the variational lower bound (VLB). Similarly, Bao et al. (2022a) explore learning a state-dependent diagonal covariance directly from data, also examining its analytical form. These methods have demonstrated improved image quality with fewer denoising steps. Learning the covariance through VLB optimization (Nichol & Dhariwal, 2021) has become a widely adopted strategy in state-of-the-art image and video generative models.
@@ 2.1 Markovian Diffusion Model: DDPM
Let $q(x_{0})$ be the true data distribution, denoising diffusion probabilistic models (DDPM) (Ho et al., 2020) constructs the following Markovian diffusion process
@@
where $q(x_{t}|x_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ and $10{<}\beta_{1}{<}\cdots{<}\beta_{T}{<}$ 1 is the pre-defined noise schedule. We can also derive the following skip-step noising distribution in a closed form
@@
Using the Tweedie’s Lemma (Efron, 2011; Robbins, 1992), we can obtain the score representation of the $\mu_{t-1}(x_{t};\theta)$
@@
For the covariance $\Sigma_{t-1}(x_{t};\theta)$, two heuristic choices were proposed in the original DDPM paper: 1. $\Sigma_{t-1}(x_{t};\theta)=\beta_{t}$, which is equal to the variance of $q(x_{t}|x_{t-1})$ and 2. $\Sigma_{t-1}(x_{t};\theta)=\tilde{\beta}_{t}$ where $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ is the variance of $q(x_{t-1}|x_{t},x_{0})$. Although heuristic, Ho et al. (2020) show that when $T$ is large, both options yield similar generation quality.
@@ 2.2 Non-Markovian Diffusion Model: DDIM
$\mu_{t{-}1}=\sqrt{\bar{\alpha}_{t{-}1}}x_{0}+\frac{x_{t}{-}\sqrt{\bar{\alpha}_{t}}x_{0}}{\sqrt{1-\bar{\alpha}_{t}}}\sqrt{1{-}\bar{\alpha}_{t{-}1}{-}\sigma_{t{-}1}^{2}}.$ (6)
@@ 3 Diffusion Models with Optimal Covariance Matching
where $H(\tilde{x})\equiv\nabla_{\tilde{x}}^{2}\log q(\tilde{x})$ and $v_{m}{\sim}p(v)$ is a Rademacher random variable (Hutchinson, 1990) with entries ±$\pm 1$ and $\odot$ denotes the element-wise product. In Table 1, we compare the generation quality of DDPM using different covariance choices. The results demonstrate that generation quality improves significantly when using the Rademacher estimator to estimate the optimal covariance, compared to the heuristic choices of $\beta$ and $\tilde{\beta}$. However, as shown in Figure 1, achieving a desirable approximation on the CIFAR10 dataset necessitates $M\geq 100$ Rademacher samples. Each calculation of $v_{m}\odot H(\tilde{x})v_{m}$ requires two network passes, resulting in a total of $2M$ network evaluations. This significantly slows down the generation speed, making it impractical for diffusion models. Inspired by Nichol & Dhariwal (2021); Bao et al. (2022a) and also the amortization technique used in variational inference (Kingma & Welling, 2013; Dayan et al., 1995), we propose to use a network to match the diagonal Hessian, which only requires one network pass to predict the diagonal Hessian in the generation time and can be done in parallel with the score/mean predictions with no extra time cost. In the next section, we introduce a novel unbiased objective to learn the diagonal Hessian from the learned score, which improves the covariance estimation accuracy and leads to better generation quality and higher likelihood estimations.
@@ 3.1 Unbiased Optimal Covariance Matching
Theorem 2 (Validity of the OCM objetcive).
The objective in Equation 11 upper bounded the grounded objective. Moreover, it attains optimal when $h_{\phi}(\tilde{x})=\text{diag}(H(\tilde{x}))$ for all $\tilde{x}\sim q(\tilde{x})$.
See the Section A.2 for a proof. The integration over $v$ in Equation 11 can be unbiasedly approximated by the Monte-Carlo integration given $M$ Rademacher samples $v_{m}\sim p(v)$. In practice, we found $M=1$ works well, which also shows the training efficiency of the proposed OCM objective. The learned $h_{\phi}$ can form the covariance approximation
@@
Algorithm 1 OCM-DDPM
@@
Algorithm 2 OCM-DDIM
@@
Figure 2: The sampling procedure OCM-DDPM and OCM-DDIM. We only demonstrate one step from $t\rightarrow t^{\prime}$ for simplicity.
@@ 3.2 Diffusion Models Applications
Given that we have access to a learned score function $\nabla_{x_{t}}\log p_{\theta}(x_{t})$, which can be obtained from any pre-trained diffusion model, we denote the Jacobian of the score by $H_{t}(x_{t})$. Assuming $M=1$ in the OCM training objective, the covariance learning objective for diffusion models can then be expressed as follows
@@
$x_{t^{\prime}}$ $=\sqrt{\bar{\alpha}}_{t^{\prime}}x_{0}+\sqrt{1-\bar{\alpha}_{t^{\prime}}}\cdot\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}x_{0}}{\sqrt{1-\bar{\alpha}_{t}}},$ (17)
@@
The sample $x_{0}\sim p_{\theta,\phi}(x_{0}|x_{t})$ can be obtained by $x_{0}=\mu_{0}(x_{t};\theta)+\epsilon\Sigma^{1/2}_{0}(x_{t};\phi)$. In the next section, we give an overview of other covariance estimation methods and compare their performance in practical examples.
@@ 3.3 Details of Training and Inference
$s_{\theta}(x_{t})$ $=\mathrm{NN}_{1}(\mathrm{BaseNet}(x_{t},t;\theta_{1});\theta_{2})$ (20)
$h_{\phi}(x_{t})$ $=\mathrm{NN}_{2}(\mathrm{BaseNet}(x_{t},t;\theta_{1});\phi)$ (21)
where $\mathrm{BaseNet}$ represents the commonly used architecture in diffusion models, such as UNet and DiT (Ho et al., 2020; Peebles & Xie, 2023). This parameterization approach only requires an additional small neural network, $\mathrm{NN}_{2}$, resulting in negligible extra computational and memory costs compared to the original diffusion models (see Section B.1 for more details). In our experiment, we fix the parameter $\theta=\{\theta_{1},\theta_{2}\}$ and train the Hessian prediction network exclusively with the proposed OCM objective. After training, samples can be generated using the algorithms summarized in Figure 2.
@@ 4 Related Covariance Estimation Methods
where $\bar{\beta}_{t^{\prime}:t}=1-\bar{\alpha}_{t^{\prime}:t}$. In optimal, $\Sigma_{t^{\prime}}(x_{t};\psi)$ in Equation 24 will recover the true covariance in Equation 8. We demonstrate the equivalence between COM-DDPM and SN-DDPM in Section A.3. However, due to the appearance of the quadratic noise term in Equation 23, SN-DDPM tends to amplify the estimation error as $t\rightarrow 0$, leading to suboptimal solutions. To mitigate this issue, Bao et al. (2022a) also propose NPR-DDPM, which models the noise prediction residual instead. We recommend referring to their paper for detailed explanations.
@@
(a) Cov Error with True Score
(b) Cov Error with Learned Score
(c) DDPM MMD v.s. Steps
(d) DDIM MMD v.s. Steps
Figure 3: Comparisons of different covariance estimation methods based on estimation error and sample generation quality. Figures (a) and (b) show the mean square error of the estimated diagonal covariance under the assumptions: (a) access to the true score, and (b) learned score of the data distribution at various noise levels. Figures (c) and (d) present the MMD evaluation against the total sampling steps in the DDPM (c) and DDIM (d) settings. We can find the proposed OCM method can achieve the lowest estimation error and consistently outperform other methods when fewer generation steps are applied.
@@ 5.1 Toy Demonstration
We first demonstrate the effectiveness of our method by considering the data distribution as a two-dimensional mixture of nine Gaussians (MoG) with means located at $[-3,0,3]\otimes[-3,0,3]$ and a standard deviation of $\sigma=0.1$, where $\otimes$ denotes Cartesian product. In this case, both the true score $\nabla_{x_{t}}\log p(x_{t})$ and the optimal covariance $\Sigma_{t^{\prime}}(x_{t})$ are available, allowing us to compare the covariance estimation error. All methods, including I-DDPM, A-DDPM, SN-DDPM, and our OCM-DDPM, assume accurate score estimation. We first learn the covariance using the true scores. Specifically, we train the covariance networks for these methods over 50,000 iterations with a learning rate of $0.001$ using an Adam optimizer (Kingma & Ba, 2014). Figure 3(a) shows the $L_{2}$ error of the estimated diagonal covariance $\Sigma_{t{-}1}(x_{t})$, and our method, OCM-DDPM, consistently achieves the lowest error compared to the other methods. In practice, the true score is not accessible. Therefore, we also conduct a comparison using a score function learned with DSM. We then apply different covariance estimation methods under the same settings. In Figure 3(b), we plot the same $L_{2}$ error for the learned score setting and find that our method achieves the lowest error for all $t$ values.
We further use the learned score and the covariance estimated by different methods to conduct DDPM and DDIM sampling for this MoG problem. For evaluation, we employ the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012), which utilizes five kernels with bandwidths $\{2^{-2},2^{-1},2^{0},2^{1},2^{2}\}$. In Figures 3(c) and 3(d), we show the MMD comparison among different methods. Specifically, we choose to conduct different diffusion steps with the skip-step scheme as discussed in Section 3.2. We find that our method consistently outperforms the others when the total time step is small, demonstrating the importance of accurate covariance estimation in the context of diffusion acceleration.
@@ 5.2 Image Modelling with Diffusion Models
Notably, while SN-DDPM often demonstrates slightly better FID performance in certain cases, it struggles with likelihood estimation, as shown in Table 3. Conversely, as discussed in the previous section, NPR-DDPM sometimes achieves superior likelihood results in Table 3, but its image generation quality is significantly worse than OCM-DDPM. Therefore, the proposed OCM-DDPM provides a more balanced model, offering a better trade-off between generation quality and likelihood estimation.
@@ 5.3 Image Modelling with Latent Diffusion Models
It is important to note that DiT was initially trained using the Improved DDPM (I-DDPM) algorithm (Nichol & Dhariwal, 2021). Therefore, the results of I-DDPM reflect the original performance of the pre-trained DiT model [^1] We use the DiT XL/2 checkpoint released in https://github.com/facebookresearch/DiT. provided by Peebles & Xie (2023). For DDPM and DDIM sampling, we discard the learned covariance in DiT and use only the learned noise prediction neural network for the posterior mean. In contrast, our OCM-DPM retains the same mean function while learning the covariance using the optimal covariance matching objective. We then generate samples with 10 sampling steps and report FID and Recall scores across different CFG coefficients on the ImageNet 256x256 dataset. The results are displayed in Figure 4, with DDPM-$\beta$ excluded due to its poor performance. It shows that our OCM-DPM method achieves the best FID performance with $\text{CFG}=2.0$. Although DDPM-$\tilde{\beta}$ and DDIM perform better in terms of FID at $\text{CFG}=1.5$, their Recall scores are lower than those of the proposed OCM methods. This discrepancy arises partly because DDPM and DDIM model
the inverse process with deterministic variance, whereas OCM methods explicitly learn the variance from data, allowing for more accurate density estimation and generating more diverse results. As Peebles & Xie (2023) report that DiT achieves the best FID performance with $\text{CFG}=1.5$, we further evaluate performance with different sampling steps under $\text{CFG}=1.5$, as shown in Table 12. The results indicate that our OCM-DPM methods offer a great improvement in sample quality and diversity.

### v4

# Improving Probabilistic Diffusion Models With Optimal Diagonal Covariance Matching

@@ Abstract
The probabilistic diffusion model has become highly effective across various domains. Typically, sampling from a diffusion model involves using a denoising distribution characterized by a Gaussian with a learned mean and either fixed or learned covariances. In this paper, we leverage the recently proposed covariance moment matching technique and introduce a novel method for learning the diagonal covariance. Unlike traditional data-driven diagonal covariance approximation approaches, our method involves directly regressing the optimal diagonal analytic covariance using a new, unbiased objective named Optimal Covariance Matching (OCM). This approach can significantly reduce the approximation error in covariance prediction. We demonstrate how our method can substantially enhance the sampling efficiency, recall rate and likelihood of commonly used diffusion models.
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019) have achieved remarkable success in modeling complex data across various domains (Rombach et al., 2022; Li et al., 2022; Poole et al., 2022; Ho et al., 2022; Hoogeboom et al., 2022; Liu et al., 2023). A conventional diffusion model operates in two stages: a forward noising process, indexed by $t\in[1,T]$, which progressively corrupts the data distribution into a Gaussian distribution via Gaussian convolution, and a reverse denoising process, which generates images by gradually transforming Gaussian noise back into coherent data samples. In traditional diffusion models, the generation process typically predicts only the mean of the denoising distribution while using a fixed, pre-defined variance (Ho et al., 2020). This approach often requires a very large number of steps, $T$, to produce high-quality, diverse samples or to achieve reasonable model likelihoods, leading to inefficiencies during inference.
To address this inefficiency, several works have proposed estimating the diagonal covariance of denoising distributions rather than relying on pre-defined variance values. For instance, Bao et al. (2022b) introduces an analytical form of isotropic, state-independent covariance that can be estimated from the data. This analytical solution achieves an optimal form under the constraints of isotropy and state independence. A more flexible approach involves learning a state-dependent diagonal covariance. Nichol & Dhariwal (2021) propose learning this form by optimizing the variational lower bound (VLB). Bao et al. (2022a) explore learning a state-dependent diagonal covariance directly from data, also examining its analytical form. These methods have demonstrated improved image quality with fewer denoising steps. Learning the covariance through VLB optimization (Nichol & Dhariwal, 2021) has become a widely adopted strategy in state-of-the-art image and video generative models.
@@ 2.1 Markovian Diffusion Model: DDPM
Let $q(x_{0})$ be the true data distribution, denoising diffusion probabilistic models (DDPM) (Ho et al., 2020) constructs the following Markovian process
@@
where $q(x_{t}|x_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ and $\beta_{1:T}$ is the pre-defined noise schedule. We can also derive the following skip-step noising distribution in a closed form
@@
With Tweedie’s Lemma (Efron, 2011; Robbins, 1992), we can have the following score representation
@@
For the covariance $\Sigma_{t-1}(x_{t};\theta)$, two heuristic choices were proposed in the original DDPM paper: 1. $\Sigma_{t-1}(x_{t};\theta)=\beta_{t}$, which is equal to the variance of $q(x_{t}|x_{t-1})$ and 2. $\Sigma_{t-1}(x_{t};\theta)=\tilde{\beta}_{t}$ where $\tilde{\beta}_{t}=(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})\beta_{t}$ is the variance of $q(x_{t-1}|x_{t},x_{0})$. Although heuristic, Ho et al. (2020) show that when $T$ is large, both options yield similar generation quality.
@@ 2.2 Non-Markovian Diffusion Model: DDIM
$\mu_{t{-}1}=\sqrt{\bar{\alpha}_{t{-}1}}x_{0}+\sqrt{1{-}\bar{\alpha}_{t{-}1}{-}\sigma_{t{-}1}^{2}}(x_{t}{-}\sqrt{\bar{\alpha}_{t}}x_{0})/\sqrt{1-\bar{\alpha}_{t}}.$ (6)
@@ 3 Diffusion Models with Optimal Covariance Matching
where $H(\tilde{x})\equiv\nabla_{\tilde{x}}^{2}\log q(\tilde{x})$ and $v_{m}{\sim}p(v)$ is a Rademacher random variable (Hutchinson, 1990) with entries ±$\pm 1$ and $\odot$ denotes the element-wise product. In Table 1, we compare the generation quality of DDPM using different covariance choices. The results demonstrate that generation quality improves significantly when using the Rademacher estimator to estimate the optimal covariance, compared to the heuristic choices of $\beta$ and $\tilde{\beta}$. However, as shown in Figure 5 in the Appendix, achieving a desirable approximation on the CIFAR10 dataset necessitates $M\geq 100$ Rademacher samples. Each calculation of $v_{m}\odot H(\tilde{x})v_{m}$ requires a forward pass and a backward propagation, leading to roughly $2M$ network evaluations in total. This significantly slows down the generation speed, making it impractical for diffusion models. Inspired by Nichol & Dhariwal (2021); Bao et al. (2022a) and also the amortization technique used in variational inference (Kingma & Welling, 2013; Dayan et al., 1995), we propose to use a network to match the diagonal Hessian, which only requires one network pass to predict the diagonal Hessian in the generation time and can be done in parallel with the score/mean predictions with no extra time cost. In the next section, we introduce a novel unbiased objective to learn the diagonal Hessian from the learned score, which improves the covariance estimation accuracy and leads to better generation quality and higher likelihood estimations.
@@ 3.1 Unbiased Optimal Covariance Matching
Theorem 2 (Validity of the OCM objective).
The objective in Equation 11 upper bounded the base objective (i.e., Equation 10 with $M\rightarrow\infty$). Moreover, it attains optimal when $h_{\phi}(\tilde{x})=\text{diag}(H(\tilde{x}))$ for all $\tilde{x}\sim q(\tilde{x})$.
See the Section A.2 for a proof. The integration over $v$ in Equation 11 can be unbiasedly approximated by the Monte-Carlo integration given $M$ Rademacher samples $v_{m}\sim p(v)$. In practice, we found $M=1$ works well (see Table 11 in the appendix for the ablation study on varying $M$ values), which also shows the training efficiency of the proposed OCM objective. The learned $h_{\phi}$ can form the covariance approximation
@@
Algorithm 1 Sampling procedure from $t\rightarrow t^{\prime}$ in OCM-DDPM
@@
Algorithm 2 Sampling procedure from $t\rightarrow t^{\prime}$ in OCM-DDIM
@@ 3.2 Diffusion Models Applications
Given access to a learned score function, $\nabla_{x_{t}}\log p_{\theta}(x_{t})$, from any pre-trained diffusion model, we denote the Jacobian of the score as $H_{t}(x_{t})$. Assuming $M=1$ in the OCM training objective, the covariance learning objective for diffusion models can be expressed as follows:
@@
$x_{t^{\prime}}$ $=\sqrt{\bar{\alpha}}_{t^{\prime}}x_{0}+\sqrt{1-\bar{\alpha}_{t^{\prime}}}/\sqrt{1-\bar{\alpha}_{t}}\cdot(x_{t}-\sqrt{\bar{\alpha}_{t}}x_{0}),$ (17)
@@
The sample $x_{0}\sim p_{\theta,\phi}(x_{0}|x_{t})$ can be obtained by $x_{0}=\mu_{0}(x_{t};\theta)+\epsilon\Sigma^{1/2}_{0}(x_{t};\phi)$. In the next section, we will compare the proposed method to other covariance estimation methods in practical examples.
@@ 3.3 Details of Training and Inference
$s_{\theta}(x_{t})=\mathrm{NN}_{1}(\mathrm{BaseNet}(x_{t},t;\theta_{1});\theta_{2}),\quad h_{\phi}(x_{t})=\mathrm{NN}_{2}(\mathrm{BaseNet}(x_{t},t;\theta_{1});\phi)$ (20)
where $\mathrm{BaseNet}$ represents the commonly used architecture in diffusion models, such as UNet and DiT (Ho et al., 2020; Peebles & Xie, 2023). This parameterization approach only requires an additional small neural network, $\mathrm{NN}_{2}$, resulting in negligible extra computational and memory costs compared to the original diffusion models (see Section B.1 for more details). In our experiment, we fix the parameter $\theta=\{\theta_{1},\theta_{2}\}$ and train the Hessian prediction network exclusively with the proposed OCM objective. After training, samples can be generated using Algorithms 1 and 2.
@@ 4 Related Covariance Estimation Methods
where $\bar{\beta}_{t^{\prime}:t}=1-\bar{\alpha}_{t^{\prime}:t}$. In optimal, $\Sigma_{t^{\prime}}(x_{t};\psi)$ in Equation 23 will recover the true covariance in Equation 8. We demonstrate the equivalence between OCM-DDPM and SN-DDPM in Section A.3. However, due to the appearance of the quadratic term in Equation 23, SN-DDPM tends to amplify the estimation error as $t\rightarrow 0$, leading to suboptimal solutions. To mitigate this issue, Bao et al. (2022a) also propose NPR-DDPM, which models the noise prediction residual instead. We recommend referring to their paper for detailed explanations.
@@ 5.1 Toy Demonstration
We first demonstrate the effectiveness of our method by considering the data distribution as a two-dimensional mixture of forty Gaussians (MoG) with means uniformly distributed over $[-40,40]\otimes[-40,40]$ and a standard deviation of $\sigma=\sqrt{40}$, where $\otimes$ denotes Cartesian product (see Figure 1(a) for visualization). In this case, both the true score $\nabla_{x_{t}}\log p(x_{t})$ and the optimal covariance $\Sigma_{t^{\prime}}(x_{t})$ are available, allowing us to compare the covariance estimation error. We then learn the covariance using the true scores by different methods to conduct DDPM and DDIM sampling for this MoG problem. For evaluation, we employ the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012), which utilizes five kernels with bandwidths $\{2^{-2},2^{-1},2^{0},2^{1},2^{2}\}$. In Figures 4(c) and 4(d), we show the MMD comparison among different methods. Specifically, we choose to conduct different diffusion steps with the skip-step scheme as discussed in Section 3.2.
(a) Training Data and Density
(b) DDPM MMD v.s. Steps
(c) DDIM MMD v.s. Steps
Figure 1: Comparisons of different covariance estimation methods. Figure (a) demonstrates the training data and the ground truth density. Figures (b) and (c) present the MMD evaluation against the total sampling steps in the DDPM (b) and DDIM (c) settings.
The results, shown in Figures 1(b) and 1(c), demonstrate that the proposed method outperforms other covariance learning approaches. Additionally, we include two methods utilizing the true diagonal and full covariance, serving as benchmarks for the best achievable performance. Notably, in this case, using the full covariance yields better performance compared to the diagonal covariance. This observation highlights the importance of accurate covariance approximation, suggesting that improved methods for learning the off-diagonal terms could lead to even better results. For further analysis, we present an additional toy experiment in Section C.1 to demonstrate the efficacy of our approach.
Figure 2: The results of FID v.s. NLL for different methods with varying numbers of sampling steps on CIFAR10 (CS). Our method consistently achieves the best trade-off between FID and NLL.
@@ 5.2 Image Modelling with Diffusion Models
Notably, while SN-DDPM often demonstrates slightly better FID performance in certain cases, it struggles with likelihood estimation, as shown in Table 3. Conversely, as discussed in the previous section, NPR-DDPM sometimes achieves superior likelihood results in Table 3, but its image generation quality is significantly worse than OCM-DDPM. Therefore, the proposed OCM-DDPM provides a more balanced model, offering a better trade-off between generation quality and likelihood estimation. To better highlight this benefit, we include a visualization in Figure 2.
@@ 5.3 Image Modelling with Latent Diffusion Models
It is important to note that DiT was initially trained using the Improved DDPM (I-DDPM) algorithm (Nichol & Dhariwal, 2021). Therefore, the results of I-DDPM reflect the original performance of the pre-trained DiT model [^1] We use the DiT XL/2 checkpoint released in https://github.com/facebookresearch/DiT. provided by Peebles & Xie (2023). For DDPM and DDIM sampling, we discard the learned covariance in DiT and use only the learned noise prediction neural network for the posterior mean. In contrast, our OCM-DPM retains the same mean function while learning the covariance using the optimal covariance matching objective. We then generate samples with 10 sampling steps and report FID and Recall scores across different CFG coefficients on the ImageNet 256x256 dataset. The results are displayed in Figure 3, with DDPM-$\beta$ excluded due to its poor performance. It shows that our OCM-DPM method achieves the best FID performance with $\text{CFG}=2.0$. Although DDPM-$\tilde{\beta}$ and DDIM perform better in terms of FID at $\text{CFG}=1.5$, their Recall scores are lower than those of the proposed OCM methods. This discrepancy arises partly because DDPM and DDIM model the inverse process with deterministic variance, whereas OCM methods explicitly learn the variance from data, allowing for more accurate density estimation and generating more diverse results. As Peebles & Xie (2023) report that DiT achieves the best FID performance with $\text{CFG}=1.5$, we further evaluate performance with different sampling steps under $\text{CFG}=1.5$, as shown in Table 13. The results indicate that our OCM-DPM methods offer a great improvement in sample quality and diversity.
@@
6 Related Work and Future Directions
In this paper, we show that improving diagonal covariance estimation can significantly enhance the performance of diffusion models. This raises a natural question: can more flexible covariance structures further improve these models? In Figure 1, we demonstrate that a full covariance structure achieves better generation quality with fewer NFEs in a toy 2D problem. However, for high-dimensional problems, the quadratic growth in parameter scale makes a full covariance approach computationally impractical. To address this, low-rank or block-diagonal covariance approximations offer promising alternatives. Developing effective training objectives for flexible covariance structures remains a compelling direction for future research.
In addition to the covariance estimation methods discussed in Section 4, there are other approaches to accelerate sampling in diffusion models. One approach involves using faster numerical solvers for differential equations with continuous timesteps (Jolicoeur-Martineau et al., 2021; Liu et al., 2022; Lu et al., 2022). Another strategy, inspired by Schrödinger bridge (Wang et al., 2021; De Bortoli et al., 2021), is to introduce a nonlinear, trainable forward diffusion process. Additionally, replacing Gaussian modelling of $p_{\theta}(x_{0}|x_{t})$ with more expressive alternatives, such as GANs (Xiao et al., 2021; Wang et al., 2022), distributional models (Bortoli et al., 2025), latent variable models (Yu et al., 2024), or energy-based models (Xu et al., 2024), can also accelerate the sampling of diffusion models.
Recently, distillation techniques have gained popularity, achieving state-of-the-art in one-step generation (Zhou et al., 2024). There are two prominent types of distillation methods. The first is trajectory distillation (Salimans & Ho, 2022; Berthelot et al., 2023; Song et al., 2023; Heek et al., 2024; Kim et al., 2023; Li & He, 2024), which focuses on accelerating the process of solving differential equations. The second type involves distillation techniques that utilize a one-step implicit latent variable model as the student model (Luo et al., 2024; Salimans et al., 2024; Xie et al., 2024; Zhou et al., 2024; Zhang et al., 2025) and distills the diffusion process into the student model by minimizing the spread divergence family (Zhang et al., 2020; 2019) through score estimation (Poole et al., 2022; Wang et al., 2024).
Beyond image modelling, our method can be straightforwardly applied to accelerate large-scale video diffusion models (Blattmann et al., 2023; Chen et al., 2023), which are based on latent diffusion models. A recent study (Zhao et al., 2024) shows that covariance estimation is crucial in mitigating the image-leakage issue in image-to-video generation problems, presenting another promising application of our method. Moreover, our method can be applied to solve inverse problems, where the optimal covariance can enhance the accuracy of posterior sampling (Chung et al., 2022; Rozet et al., 2024). While this work focuses on the DDPM and DDIM sampler, incorporating the improved covariance into the general stochastic and deterministic differential solver (Song et al., 2021; Karras et al., 2022) also represents exciting directions for future research.